Monitoring - We are continuing to monitor this situation for any further issues at this time.
Jul 26, 19:37 UTC
Investigating - We are aware of a networking issue that is affecting a small portion of customers in our Singapore location. We are working on a resolution and will post an update as soon as we have more information. For the meantime, we have found that rebooting your Linode will partially resolve the issue.
Jul 26, 08:50 UTC
Linode Manager and API Operational
London Operational
Newark Operational
Atlanta Operational
Dallas Operational
Fremont Operational
Tokyo Operational
Singapore Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Past Incidents
Jul 28, 2015

No incidents reported today.

Jul 27, 2015

No incidents reported.

Jul 25, 2015

No incidents reported.

Jul 24, 2015

No incidents reported.

Jul 23, 2015
Resolved - We have mitigated a large-scale DDoS attack and the packet loss has subsided. We are continuing to monitor traffic to ensure this attack does not resume.
Jul 23, 17:11 UTC
Identified - This packet loss is due to a large-scale DDoS attack, which we are currently working on mitigating.
Jul 23, 16:04 UTC
Resolved - Our Backup Administrators have completed maintenance in Fremont and Dallas. We thank you for your patience.
Jul 23, 12:39 UTC
Identified - The issue has been identified and a fix is being implemented.
Jul 20, 20:12 UTC
Investigating - Our Administrators are performing emergency maintenance on the Backups Service for a subset of customers in Fremont and Dallas. We are working quickly to restore normal Backups Service for affected customers.
Jul 20, 20:12 UTC
Jul 22, 2015
We experienced a hardware failure on some network equipment in our Fremont data center, and as a result, some customer traffic was affected for a brief period. The issue is now resolved and you should see normal connectivity in Fremont at this time.
Jul 22, 13:11 UTC
Jul 21, 2015
Completed - The scheduled maintenance has been completed.
Jul 21, 18:45 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 21, 17:45 UTC
Scheduled - We will be performing network maintenance on the date and time listed above. We do not expect any downtime during this maintenance, however a brief period of packet loss or increased latency may be observed.
Jul 21, 17:34 UTC
Completed - Scheduled maintenance on the Linode Manager and API has been completed. All systems should be fully operational at this time.
Jul 21, 02:07 UTC
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 21, 02:00 UTC
Scheduled - A maintenance window has been scheduled for the Linode Manager and API on Monday, July 20th between 10:00pm and 11:00pm EDT (GMT-4). We expect this maintenance to take approximately 30 minutes, but it is possible it will last the full 1 hour duration of the window. Linodes will not be affected by this maintenance, but the Linode Manager and API will be unavailable.
Jul 20, 18:51 UTC
Jul 19, 2015

No incidents reported.

Jul 18, 2015

No incidents reported.

Jul 17, 2015

No incidents reported.

Jul 16, 2015
We are aware of an issue that affected IPv6 connectivity in our Fremont facility between 11:45p and 12:45a EDT. We will update again when we recieve more information from our upstream provider.
Jul 16, 06:09 UTC
Resolved - At this time, the issue with Linode's repository mirrors has been resolved. If you are experiencing any lingering issues, please contact Linode Support.
Jul 16, 04:14 UTC
Identified - We've identified an issue with Linode's repository mirrors, which may prevent the installation of packages from our mirrors. This issue is affecting Linode's repository mirrors in multiple datacenters. Our Administrators are working to resolve this issue as quickly as possible.
Jul 15, 14:06 UTC
Jul 14, 2015

No incidents reported.